OpenStudy (anonymous):

how do you get the answer in here using implicit differentiation? x^2 + y^2 = r^2

OpenStudy (somethingawesome):

I take it we want to find dy/dx? To use implicit differentiation, we take the derivative of each term as usual, keeping in mind that y is a function of x, so the chain rule makes the dy/dx term appear. Taking the derivative of both sides then gets \[2x + 2y \frac{dy}{dx} = 0\]since r is a constant. Now we just have to solve for dy/dx: \[ \frac{dy}{dx} = \frac{-x}{y}\]

OpenStudy (anonymous):

so if there's a constant in the formula, it would be zero automatically then?

OpenStudy (somethingawesome):

The derivative of a constant is still zero, yes.